# Constraining electromagnetic couplings of ultralight scalars from compact stars

### v1

@@ IV EM wave propagation in the background of a long-range scalar field
$\delta z=\frac{\Delta k}{k}\sim 10^{-4}\Big{(}\frac{g_{\phi\gamma\gamma}}{10^{-16}~{}\mathrm{GeV}^{-1}}\Big{)}^{2}\Big{(}\frac{2.1~{}\mathrm{GHz}}{\omega}\Big{)}\Big{(}\frac{B_{0}}{3.93\times 10^{16}~{}\mathrm{G}}\Big{)}^{2}\Big{(}\frac{R}{10~{}\mathrm{km}}\Big{)}^{4}\Big{(}\frac{1.4~{}M_{\odot}}{M}\Big{)}^{3},$ (22)
where we have used Eqs. 8 and 9 to describe the scalar field and $k\approx\omega$ at the leading order. If the redshift measurements have a precision of $\sim 10^{-4}$ Rowlinson:2010jb , we would get sensitivity of $g_{\phi\gamma\gamma}\sim 10^{-16}$ GeV-1 to the scalar-photon coupling.
With the advancements in precision atomic clocks, it may be possible in future to determine the wavelength (or frequency) of a particular spectral line emitted by the magnetar to a precision of $\Delta k/k\sim 10^{-18}$. From Eq. 22, this precision would correspond to the sensitivity in $g_{\phi\gamma\gamma}$ of $\sim 10^{-23}$ GeV-1. This prospective bound has been indicated by a dashed purple line in FIG. 1. The shift in the wavelength would be more pronounced for compact stars that have strong surface magnetic fields, larger dimensions, and emit signals detectable at lower frequencies.
@@ V Scalar field radiation from an isolated compact star
In the situations considered in previous sections, the scalar field was coupled to a static source, resulting in a long-range $1/r$ scalar field profile outside the magnetized star. Since a static source does not emit scalar radiation, no scalar field radiation originates from the star in these situations. To investigate the impact of scalar radiation on pulsar spin-down, we now consider a scenario where the EM fields of the magnetar are oscillating with time. We consider a skewed rotator model, where the magnetic moment axis of the star makes an angle $\alpha$ with its rotation axis. In this model, the magnetic and electric fields at any space-time point can be written as cite-key
$\begin{split}\textbf{B}=\frac{B_{0}R^{3}}{2r^{3}}\Big{[}(3\cos\theta_{m}\sin\theta\cos\varphi-\sin\alpha\cos\Omega t)(\sin\theta\cos\varphi\,\hat{r}\,+\cos\theta\cos\varphi\,\hat{\theta}-\sin\varphi\,\hat{\varphi})\,+\\ (3\cos\theta_{m}\sin\theta\sin\varphi-\sin\alpha\sin\Omega t)(\sin\theta\sin\varphi\,\hat{r}+\cos\theta\sin\varphi\,\hat{\theta}+\cos\varphi\,\hat{\varphi})+\\ (3\cos\theta_{m}\cos\theta-\cos\alpha)\times(\cos\theta\,\hat{r}-sin\theta\,\hat{\theta})\Big{]},\end{split}$ (23)
and
$\begin{split}\textbf{E}=-\frac{B_{0}\Omega R^{5}}{2r^{4}}\Big{[}\{\cos\alpha(3\cos^{2}\theta-1)+3\sin\alpha\sin\theta\cos\theta\cos(\Omega t-\varphi)\}\hat{r}\,+\\ 2\{\cos\alpha\sin\theta\cos\theta+\sin\alpha\sin^{2}\theta\cos(\Omega t-\varphi)\}\hat{\theta}\Big{]},\end{split}$ (24)
where the magnetic colatitude $\theta_{m}$ is the angle between the magnetic moment axis and the line of sight. It is expressed as $\cos\theta_{m}=\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos(\Omega t-\varphi)$, where $\alpha$ is the angle between the rotational axis and the magnetic moment axis, and $\theta$ is the angle between the rotational axis and the line of sight. In the limit $\alpha\rightarrow 0$, Eqs. 23 and 24 reduce to Eqs. 1 and 2, respectively. Thus, for radiation, one needs a non-zero $\alpha$.
The source charge density for $\phi$ may be written as $\rho_{\phi}(\mathbf{r},t)=g_{\phi\gamma\gamma}(\mathbf{B}^{2}-\mathbf{E}^{2})$, where Eqs. 23 and 24 give
$\mathbf{B}^{2}-\mathbf{E}^{2}\approx-\frac{3}{2}\frac{B^{2}_{0}R^{6}}{r^{6}}\cos\theta_{m}\sin\alpha\sin\theta\cos(\Omega t-\varphi).$ (25)
Here we omit terms involving $\Omega R\ll 1$ and higher-order terms of $\mathcal{O}(\alpha^{2})$ for small $\alpha$. The dominant contribution to the radiation comes from the fundamental harmonic, since the power radiated in the higher harmonics is suppressed by the powers of the velocity of the rotating star, and hence can be neglected. We also remove the time-independent terms which do not contribute to the radiation.
We use the far-field and long-wavelength approximation. Following Krause:1994ar and taking the time-average over the rotation period, we obtain the rate of scalar dipole radiation as
@@
where $k^{2}=\Omega^{2}-m^{2}_{\phi}$ and $dS_{n}$ is the solid angle. The time-averaged dipole moment $\mathbf{p}_{\Omega}$ is given as
[…]
where $P=2\pi/\Omega$ is the rotational time period. Therefore, Using Eqs. 25, 26 and 27, we obtain
$\frac{dE}{dt}\approx\frac{\pi}{48}g^{2}_{\phi\gamma\gamma}B^{4}_{0}R^{8}\Omega^{4}\sin^{2}(2\theta_{m})\Big{(}1-\frac{m^{2}_{\phi}}{\Omega^{2}}\Big{)}^{3/2},$ (28)
which contributes to the pulsar spin-down, provided $m_{\phi}<\Omega$.
@@ VI Constraints from Observations
On the other hand, there are bounds that do not need the scalar to be the DM. The astrophysical bounds from globular clusters Dolan:2022kul are obtained by calculating the ratio of the energy losses due to the scalar in the asymptotic giant branch to the horizontal branch stars. The bounds from the Eöt-Wash experiment Hees:2018fpg , fifth force experiments Fischbach:1996eq ; Adelberger:2003zx ; Konopliv2011 ; Fienga:2023ocw and MICROSCOPE experiment Berge:2017ovy are obtained from the precision tests of Einstein’s equivalence principle, using laboratory measurements or astrophysical observations. These bounds are relevant for comparison and complementarity with our work.
@@ VI.1 Search for a new long-range force in a double pulsar binary
Figure 1: Bounds on $g_{\phi\gamma\gamma}$ derived from the measurements of the electromagnetic radiation by a scalar-induced magnetic field (blue shaded region), and pulsar spin-down caused by scalar radiation (red shaded region). The prospective bound from possible future constraints on a new long-range force from a pulsar binary pair with large surface magnetic fields $B_{0}$ is shown as a magenta dashed line, while the prospective bound from the measurement of the photon wavenumber using a future precision atomic clock is shown as a purple dashed line. Existing constraints are shown as gray-shaded areas.
@@
However, since the gravitational and the long-range force have the same spatial dependence, it would be impossible to separate their contributions by simply measuring the attractive force between them. Indeed, a change in the magnitude of the force would be mimicked by a change in the measured values of the masses. This no-go situation may be circumvented if the values of the masses are determined independently by some means other than the gravitational measurements, or we have access to a third body that is gravitationally bound with the binary but has a magnetic field much different than the two compact stars.
@@ VI.2 Electromagnetic radiation due to the scalar-induced magnetic field
$\dot{P}$ $4.20\times 10^{-13}~{}\mathrm{s~{}s^{-1}}$ Lyne:2014qqa $115.7\times 10^{-12}~{}\mathrm{s~{}s^{-1}}$ Marsden:2001tw $1.86\times 10^{-7}~{}\mathrm{s~{}s^{-1}}$owlinson:2013ue
$B_{0}$ $8.5\times 10^{12}~{}\mathrm{G}$ Lyne:2014qqa ; Khelashvili:2024sup $9.41\times 10^{14}~{}\mathrm{G}$ Marsden:2001tw $39.26\times 10^{15}~{}\mathrm{G}$ Rowlinson:2013ue
@@
The strongest bound on the coupling is obtained from GRB 080905A, which is valid for $m_{\phi}\lesssim 2\times 10^{-11}$ eV. This bound is stronger than the astrophysical bound obtained using globular clusters, and is stronger than the current fifth-force bound for $m_{\phi}\lesssim 1.5\times 10^{-22}~{}\mathrm{eV}$. This constraint is depicted in FIG. 1 by the blue-shaded region. The bounds for Crab pulsar and SGR 1806-20 are weaker than GRB 080905A and we do not show them in the figure.
@@ VI.3 Spin-down of compact stars due to scalar radiation
To establish bounds on the scalar-photon coupling, we assume that $dE/dt$ includes contributions from both standard physics and from Eq. 28. We also take $\sin^{2}(2\theta_{m})=1$ for a conservative limit. Note that, from the kinematic factor $(1-m_{\phi}^{2}/\Omega^{2})^{3/2}$ in Eq. 28, there will be no scalar radiation for $m_{\phi}>\Omega$. Therefore, the bounds are valid only for $m_{\phi}\lesssim\Omega$. Note that this range is smaller than the range $m_{\phi}\lesssim 1/R$ in the preceding subsection, since $\Omega\sim 10^{-13}$ eV while $1/R\sim 10^{-11}$ eV.
@@
$dE/dt$ $4.5\times 10^{38}~{}\mathrm{erg/s}$Manchester:2004bp ; Palomba:1999su ; Lyne:2014qqa $1.4\times 10^{36}~{}\mathrm{erg/s}$ MEREGHETTI20111317 $1.8\times 10^{48}~{}\mathrm{erg/s}$ Rowlinson:2013ue
$g_{\phi\gamma\gamma}$ $\lesssim 6\times 10^{-14}~{}\mathrm{GeV}^{-1}$ $\lesssim 2\times 10^{-14}~{}\mathrm{GeV}^{-1}$ $\lesssim 2\times 10^{-16}~{}\mathrm{GeV}^{-1}$
@@
In FIG. 1, the constraint obtained from the spin-down of GRB 080905A is shown in the red-shaded region, while the bounds from the Crab pulsar and SGR 1806-20 are not depicted, as they are weaker in comparison. The constraints from SGR 1806-20 and GRB 080905A are stronger than the astrophysical bounds from globular clusters, and surpass those of the fifth-force bounds in the ultra-low scalar mass limit $(m_{\phi}\lesssim 4.5\times 10^{-24}~{}\mathrm{eV})$. However, they are still weaker than the bounds from Eöt-Wash Hees:2018fpg and MICROSCOPE Berge:2017ovy experiments. Note that the bounds obtained from GRB 080905A could be improved by almost four orders of magnitude for a magnetar with a magnetic field $B_{0}\sim 10^{18}~{}\mathrm{G}$, making it competitive with the fifth-force constraints over a wider range of $m_{\phi}$. More sensitive pulsar spin-down measurements could lead to bounds exceeding those set by the laboratory tests for the equivalence principle.

### v2

@@ IV EM wave propagation in the background of a long-range scalar field
$\delta z=\frac{\Delta k}{k}\sim 10^{-4}\Big(\frac{g_{\phi\gamma\gamma}}{10^{-15}~\mathrm{GeV}^{-1}}\Big)^{4}\Big(\frac{2.1~\mathrm{GHz}}{\omega}\Big)^{2}\Big(\frac{B_{0}}{3.93\times 10^{16}~\mathrm{G}}\Big)^{4}\Big(\frac{R}{10~\mathrm{km}}\Big)^{8}\Big(\frac{1.4~M_{\odot}}{M}\Big)^{6},$ (22)
where we have used Eqs. 7 and 8 to describe the scalar field, and taken $k\approx\omega$ at the leading order. If the redshift measurements have a precision of $\sim 10^{-4}$ Rowlinson:2010jb , we would get a sensitivity of $g_{\phi\gamma\gamma}\sim 10^{-15}$ GeV-1 to the scalar-photon coupling. The shift in the wavenumber would be more pronounced for compact stars that have strong surface magnetic fields, larger dimensions, and emit signals detectable at lower frequencies.
With the advancements in precision atomic clocks, it may be possible to determine the wavelength (or frequency) of a particular spectral line emitted by the magnetar to a precision of $\Delta k/k\sim 10^{-18}$. From Eq. 22, this precision would correspond to a sensitivity of $g_{\phi\gamma\gamma}\sim 10^{-19}~\mathrm{GeV}^{-1}$. This prospective bound has been indicated by a dashed purple line in FIG. 1. The sensitivity can be further enhanced by observing low-frequency photons, which are accessible to radio telescope facilities such as the Square Kilometre Array (SKA) Huege:2016jvc and LOw Frequency ARray (LOFAR) Geyer:2017lbi ; vanHaarlem2013 . The constraint on $g_{\phi\gamma\gamma}$ can be further strengthened through the use of nuclear clocks, entangled clock networks, or high-precision space-based timekeeping systems Kessler:2014sct ; Arakawa:2023gyq ; Fuchs:2024xvc ; PhysRevA.81.030302 ; Yang:2025agr , by employing a large number of atoms, extending the interrogation time, and minimizing environmental noise.
In particular, conventional atomic clocks operate at the standard quantum limit (SQL) with $N$ uncorrelated atoms, the phase evolution scales as $1/\sqrt{N}$. If the atoms are prepared in maximally entangled (fully correlated) states, the ensemble behaves as a single coherent quantum system and the precision can reach the Heisenberg limit (HL), where the phase evolution scales as $1/N$. While these conditions represent near-ideal scenarios, ongoing technological developments are steadily improving the performance of realistic systems.
The corresponding sensitivity curve for a relative wavenumber deviation of $\Delta k/k\sim 10^{-24}$ at a frequency of $\omega\sim 10~\mathrm{MHz}$ (a frequency accessible at the LOFAR telescope) is shown as a green dashed line in FIG. 1
In the regime $\omega>m_{\gamma}/2$, the imaginary part of the wavenumber, $k_{I}$, leads to a damping of the wave amplitude, indicating photon absorption in the scalar field medium. This attenuation manifests in the decay of the photon’s intensity, which is related to the electric field as $I\propto\mathbf{E}^{2}\propto E_{0}^{2}e^{-2k_{I}x}$. Consequently, if a photon with initial intensity $I_{0}$ propagates through an absorptive medium of thickness $x$, the transmitted intensity becomes $I=I_{0}e^{-\alpha x}$, where $\alpha$ represents the absorption coefficient. Within the framework of our analysis, [^6] There could be enhancement in the photon intensity if $(\nabla\phi\cdot\bm{k})$ picks negative sign.
$\alpha=2k_{I}=m_{\gamma}=\frac{1}{x}\ln\Big(\frac{I_{0}}{I}\Big).$ (23)
Therefore, unlike the photon wavelength shift—which appears only at order $\mathcal{O}(g^{4}_{\phi\gamma\gamma})$—the absorption coefficient arises at the order $\mathcal{O}(g^{2}_{\phi\gamma\gamma})$.
The scalar-induced redshift may be distinguished from the possible effects of the gravitational redshift and plasma dispersion. The gravitational redshift is independent of frequency while the scalar-induced redshift goes as $1/\omega^{2}$. Plasma dispersion effects lead to the same $1/\omega^{2}$ dependence; however, they are accompanied by Faraday rotation which is absent in the scalar-induced case.
@@ V Scalar field radiation from an isolated compact star
In the situations considered in previous sections, the scalar field was coupled to a static source, resulting in a long-range $1/r$ scalar field profile outside the magnetized star. Since a static source does not emit scalar radiation, no scalar field radiation originates from the star in these situations. To investigate the impact of scalar radiation on pulsar spin-down, we now consider a scenario where the EM fields of the magnetar are oscillating with time. We consider a skewed rotator model, where the magnetic moment axis of the star makes an angle $\alpha$ with its rotation axis. In this model, the magnetic field at any space-time point can be written as cite-key
@@
where the magnetic colatitude $\theta_{m}$ is the angle between the magnetic moment axis and the line of sight. It is expressed as $\cos\theta_{m}=\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos(\Omega t-\varphi)$, where $\alpha$ is the angle between the rotational axis and the magnetic moment axis, and $\theta$ is the angle between the rotational axis and the line of sight. In the limit $\alpha\rightarrow 0$, Eq. 24 reduces to Eq. 1. Thus, for radiation, one needs a non-zero $\alpha$.
The source charge density for $\phi$ may be written as $\rho_{\phi}(\mathbf{r},t)=g_{\phi\gamma\gamma}(\mathbf{B}^{2}-\mathbf{E}^{2})$, where Eq. 24 gives
$\mathbf{B}^{2}-\mathbf{E}^{2}\approx-\frac{3}{2}\frac{B_{0}^{2}R^{6}}{r^{6}}\Big[\sin\alpha\cos\alpha\sin\theta\cos\theta\cos(\Omega t-\phi)+\frac{1}{2}\sin^{2}\alpha\sin^{2}\theta\cos^{2}(\Omega t-\phi)\Big],$ (25)
and we omit terms involving $\Omega R\ll 1$, and hence, $\mathbf{E^{2}}$. We also remove the time-independent terms which do not contribute to the radiation.
In Eq. 25, the term $\mathbf{B}^{2}-\mathbf{E}^{2}\propto\sin\theta\cos\theta$ exhibits a quadrupolar angular dependence, while the $\sin^{2}\theta$ component contains both a time-independent DC part and a quadrupolar contribution. Consequently, the first term leads to quadrupolar radiation at the rotation frequency $\Omega$, whereas the second term produces radiation at twice the frequency, $2\Omega$, in addition to the non-radiative DC component. Therefore, unlike the pseudoscalar axion case Khelashvili:2024sup , no scalar dipole radiation arises in this scenario.
The corresponding source charge densities for radiation at $\Omega$ and $2\Omega$ can therefore be written as
$\rho_{+\Omega}(\mathbf{r})=\frac{3}{2}g_{\phi\gamma\gamma}\mu^{2}\sin(2\alpha)\sqrt{\frac{8\pi}{15}}\frac{Y_{2,-1}(\theta,\phi)}{r^{6}},~~~\rho_{+2\Omega}(\mathbf{r})=-\frac{3}{4}g_{\phi\gamma\gamma}\mu^{2}\sin^{2}\alpha\sqrt{\frac{32\pi}{15}}\frac{Y_{2,-2}(\theta,\phi)}{r^{6}},$ (26)
where $\mu=B_{0}R^{3}/2$ and $Y_{l,m}(\theta,\phi)$ are spherical harmonics.
The scalar field satisfies the wave equation
$(\nabla^{2}+k^{2})\phi_{\omega}(\mathbf{r})=-\rho_{\omega}(\mathbf{r}),~~~k=\sqrt{\omega^{2}-m^{2}_{\phi}},$ (27)
whose solution, using the retarded Green’s function, is
$\phi_{\omega}(\mathbf{r})=\int d^{3}r^{\prime}\frac{e^{ik|\mathbf{r}-\mathbf{r^{\prime}}|}}{4\pi|\mathbf{r}-\mathbf{r^{\prime}}|}\rho_{\omega}(\mathbf{r^{\prime}}).$ (28)
Defining the multipole moments as
$Q_{lm}(\omega)=\int d^{3}r^{\prime}\rho_{\omega}(\mathbf{r}^{\prime}){r^{\prime}}^{l}Y_{l,m}^{*}(\hat{r}^{\prime}),$ (29)
the scalar field in the far-field and long-wavelength limit becomes
$\phi_{\omega}(\mathbf{r})\simeq\frac{e^{ikr}}{r}\sum_{lm}i^{l}\frac{k^{l}}{(2l+1)!!}Y_{l,m}(\hat{r})Q_{lm}(\omega).$ (30)
For a real scalar field, the radial energy flux is
$T^{0r}=\dot{\phi}\partial_{r}\phi,~~~\langle S_{r}\rangle=\langle T^{0r}\rangle=\frac{\omega k}{2}|\phi_{\omega}(\mathbf{r})|^{2},$ (31)
where the angle brackets denote time-averaging over one period. The total power radiated in scalar waves is then
$P_{\omega}=\int d\Omega r^{2}\langle S_{r}\rangle=\frac{\omega k}{2}\sum_{lm}\Big(\frac{k^{l}}{(2l+1)!!}\Big)^{2}|Q_{lm}(\omega)|^{2}.$ (32)
Hence, the scalar quadrupole radiation power at frequencies $\Omega$ and $2\Omega$ are given by
$P_{\Omega}\simeq\frac{1}{80}g^{2}_{\phi\gamma\gamma}B^{4}_{0}R^{10}\Omega^{6}\Big(1-\frac{m^{2}_{\phi}}{\Omega^{2}}\Big)^{5/2}\sin^{2}2\alpha,~~~P_{2\Omega}\simeq\frac{1}{100}g^{2}_{\phi\gamma\gamma}B^{4}_{0}R^{10}\Omega^{6}\Big(1-\frac{m^{2}_{\phi}}{4\Omega^{2}}\Big)^{5/2}\sin^{4}\alpha.$ (33)
Therefore, scalar radiation at frequency $\Omega$ is allowed when $m_{\phi}\lesssim\Omega$, while radiation at $2\Omega$ occurs only if $m_{\phi}\lesssim 2\Omega$. Also, for a typical pulsar, the misalignment angle is small and $P_{\Omega}$ dominates over $P_{2\Omega}$. Scalar radiation can occur only when the pulsar’s magnetic axis is tilted with respect to its rotation axis. In the aligned-rotator limit $\alpha\to 0$, where the two axes coincide, the scalar radiation vanishes.
@@ VI Constraints from Observations
On the other hand, there are bounds that do not need the scalar to be the DM. The astrophysical bounds from globular clusters Dolan:2022kul are obtained by calculating the ratio of the energy losses due to the scalar in the asymptotic giant branch to the horizontal branch stars. The bounds from the Eöt-Wash experiment Hees:2018fpg , fifth force experiments Fischbach:1996eq ; Adelberger:2003zx ; Konopliv2011 ; Fienga:2023ocw and MICROSCOPE experiment Berge:2017ovy are obtained from the precision tests of Einstein’s equivalence principle, using laboratory measurements or observations in solar system. These bounds are relevant for comparison and complementarity with our work.
@@ VI.1 Search for a new long-range force in a double pulsar binary
Figure 1: Bounds on $g_{\phi\gamma\gamma}$ derived from the measurements of the electromagnetic radiation by a scalar-induced magnetic field (blue shaded region), and pulsar spin-down caused by scalar radiation (red shaded region). The prospective bound from possible future constraints on a new long-range force from a pulsar binary pair with large surface magnetic fields $B_{0}$ is shown as a magenta dashed line, while the prospective bound from the measurement of the photon wavenumber using a future precision atomic clock ($\Delta k/k\sim 10^{-18},\omega=2.1$ GHz) is shown as a purple dashed line. The sensitivity obtainable with $\Delta k/k\sim 10^{-24}$ at $\omega\sim 10$ MHz has also been shown with a green dashed line. The existing constraints are shown as gray-shaded areas.
@@
However, since the gravitational and the scalar-induced long-range force have the same spatial dependence to the leading order, it would be difficult to separate their contributions by simply measuring the attractive force between them. Indeed, a change in the magnitude of the force could be approximately mimicked by a change in the measured values of the masses. This quasi-no-go situation may be circumvented if the values of the masses are determined independently by some means other than the gravitational measurements, or we have access to a third body that is gravitationally bound with the binary but has a magnetic field much different than the two compact stars.
@@
Precise measurements of post-Keplerian parameters such as the mass ratio, orbital period decay, orbital inclination, periastron advance, Einstein delay, and Shapiro delay enable the determination of individual pulsar masses. Currently, pulsar mass measurements can reach accuracies of about $\mathcal{O}(10^{-3}-10^{-4})$. However, upcoming radio telescopes like MeerKAT Hu:2020ubl and SKA Huege:2016jvc , along with future space-based GW observatories like LISA Thrane:2019lwv , are expected to push this precision further to an unprecedented level of $\mathcal{O}(10^{-6}-10^{-7})$. Specifically, projected measurements of the masses of Pulsars A and B in the double pulsar system PSR J0737–3039A/B by MeerKAT are expected to reach precisions of $\mathcal{O}(10^{-5})$ and $\mathcal{O}(10^{-6})$, respectively Hu:2022jiq . Moreover, MeerKAT is anticipated to measure the moment of inertia of Pulsar A with a precision of about $11\%$. This measurement will provide important constraints on the NS EOS, thereby refining the mass estimates of the pulsars even further Hu:2020ubl .
@@ VI.2 Electromagnetic radiation due to the scalar-induced magnetic field
$\dot{P}$ $4.20\times 10^{-13}~\mathrm{s~s^{-1}}$ Lyne:2014qqa $115.7\times 10^{-12}~\mathrm{s~s^{-1}}$ Marsden:2001tw $1.86\times 10^{-7}~\mathrm{s~s^{-1}}$Rowlinson:2013ue
$B_{0}$ $(6.9-8.5)\times 10^{12}~\mathrm{G}$ Khelashvili:2024sup $(8-25)\times 10^{14}~\mathrm{G}$ MEREGHETTI20111317 $(27.1-49.5)\times 10^{15}~\mathrm{G}$ Rowlinson:2013ue
@@
In order to put bounds on $g_{\phi\gamma\gamma}$, we express the observed surface magnetic field as
$B_{\mathrm{surf}}^{2}\simeq B_{0}^{2}+B_{\phi}^{2}(g_{\phi\gamma\gamma},B_{0},M,R)\;.$ (38)
We further impose the condition that the additional magnetic contribution $B_{\phi}$ does not exceed the $1\sigma$ measurement uncertainty in $B_{0}$. Uncertainties in the stellar mass and radius are treated as nuisance parameters and propagated through $B_{\phi}$.
The strongest bound on the coupling at the polar cap is obtained from GRB 080905A, which is valid for $m_{\phi}\lesssim 2\times 10^{-11}$ eV. This bound is eight orders of magnitude stronger than the astrophysical bound obtained using globular clusters, and is stronger than the current fifth-force bound for $m_{\phi}\lesssim 1.5\times 10^{-22}~\mathrm{eV}$. This constraint is depicted in FIG. 1 by the blue-shaded region. The bounds for Crab pulsar and SGR 1806-20 are weaker than GRB 080905A and we do not show them in the figure.
The reason why our method yields stronger bounds than that from the fifth-force measurements at ultralight masses is as follows. The fifth-force experiments investigate the derivative of a generic Yukawa potential across small and large length scales to identify any deviations from standard gravity. In the very long-range limit, the mass of the mediator approaches zero, making the Yukawa potential indistinguishable from the standard Newtonian potential. In our method, the constraints on the coupling are valid as long as the scalar mass is smaller than the relevant inverse-distance scale in the observed system: the inverse of the distance between the binaries, the inverse radius of the compact star, or the spin frequency of the compact star. Therefore, our results stay valid as $m_{\phi}$ goes to zero.
@@ VI.3 Spin-down of compact stars due to scalar radiation
To constrain the scalar-photon coupling, we assume that the total energy loss rate of the pulsar receives contributions from both standard mechanisms and the scalar radiation given by Eq. 33. Since the inclination angle $\alpha$ is generally small, scalar emission at frequencies $m_{\phi}\lesssim\Omega$ is expected to dominate. For a conservative estimate, we set $\sin^{2}2\alpha=1$. It is worth noting that this mass range is narrower than the condition $m_{\phi}\lesssim 1/R$ discussed in the previous subsection, as for Crab pulsar $\Omega\sim 10^{-13}~\mathrm{eV}$ while $1/R\sim 10^{-11}~\mathrm{eV}$.
The spin-down bounds are derived by comparing the observed spin-down luminosity $\dot{E}_{\mathrm{obs}}$ (Eq. 35), with the theoretical sum of the standard electromagnetic loss $\dot{E}_{\mathrm{EM}}$ (Eq. 36) and the scalar-radiation power $\dot{E}_{\phi}$ (Eq. 33). Using the measured $P,\dot{P}$, and $B_{0}$ from the vacuum dipole model, we require $\dot{E}_{\mathrm{EM}}+\dot{E}_{\phi}\leq\dot{E}_{\mathrm{obs}}$. We obtain a bound on $g_{\phi\gamma\gamma}$ by requiring that the scalar-induced energy loss $\dot{E}_{\phi}$ remains within the $1\sigma$ measurement uncertainty of the observed electromagnetic spindown power $\dot{E}_{\mathrm{EM}}$.
@@
$dE/dt$ $(4.49-4.51)\times 10^{38}~\mathrm{erg/s}$Khelashvili:2024sup $(0.5-1.4)\times 10^{36}~\mathrm{erg/s}$ MEREGHETTI20111317 $(0.7-3.8)\times 10^{48}~\mathrm{erg/s}$ Rowlinson:2013ue
$g_{\phi\gamma\gamma}$ $\lesssim 2\times 10^{-11}~\mathrm{GeV}^{-1}$ $\lesssim 7\times 10^{-9}~\mathrm{GeV}^{-1}$ $\lesssim 2\times 10^{-14}~\mathrm{GeV}^{-1}$
@@
In FIG. 1, the constraint obtained from the spin-down of GRB 080905A is shown in the red-shaded region, while the bounds from the Crab pulsar and SGR 1806-20 are not depicted, as they are weaker in comparison. The constraints from Crab pulsar and GRB 080905A are stronger than the astrophysical bounds from globular clusters. However, they are still weaker than the bounds from Eöt-Wash Hees:2018fpg and MICROSCOPE Berge:2017ovy experiments. More sensitive pulsar spin-down measurements could lead to bounds exceeding those set by the laboratory tests for the equivalence principle.